Demo file for DVIWIN 2.9 December 1, 1994 Hippocrates Sendoukas isendo@leon.nrcps.ariadne-t.gr

This driver lets you preview and print DVI files under MS-Windows 3.1 and Windows/NT. Its main advantages are: speed, compatibility with any raster device with a Windows driver, and easy integration of text and graphics. All screen and printer handling is done through Windows, so it should work with any video card and any printer supported by Windows. You can insert arbitrary graphics files produced by most Windows applications, or other standard graphics files. This version supports WMF, BMP, PCX, MSP, GIF and XPM files; other formats (eg., TIFF, HPGL, etc.) can also be imported provided that you have the appropriate graphics filter. The following graph was made with Excel and it will come out in color if you print it on a color printer.

Sample Figure: Utility surface for two goods

The driver also understands emTEX and tpic specials. The box and the shading of this paragraph was made with a few simple tpic specials.

The program also supports certain specials that facilitate the integration of the driver with an editor. This demo file contains a few such specials; if you use PFE as your editor, you can see a demonstration of the technique by following the “Inverse Search” instructions in the “readme” file, and double–clicking the left mouse button at any point on the DVI surface.

These specials have been inserted by a couple of macros inside this file; note that these macros are not enough to generate all the required specials. Therefore, the potential of this mechanism is not fully realized. The help file contains the information about the placing of the required specials; these specials can be generated either through TEX macros, or through simple modifications to TEX. The next version of BCTEX is expected to generate these specials automatically. This will vastly improve the accuracy of the mechanism.

For more information on using the programs, please read or print the files “dviwin.wri”, “wbr.wri” and “clipmeta.wri”. If you encounter any installation or operation problems, read the file “helpme.wri” and if it does not resolve the problem, please let me know.